You want to buy a $245,000 home. You plan to pay 15% as a down payment, and take out a 30 year fixed loan for the rest.
Round all answers to the nearest cent as needed.
a) How much is the loan amount going to be?
$

b) What will your monthly payments be if the interest rate is 6.2%?
$

c) What will your monthly payments be if the interest rate is 7.2%?
$

Answer by Theo(13342)

You want to buy a $245,000 home. You plan to pay 15% as a down payment, and take out a 30 year fixed loan for the rest.
Round all answers to the nearest cent as needed.
a) How much is the loan amount going to be?
b) What will your monthly payments be if the interest rate is 6.2%?
c) What will your monthly payments be if the interest rate is 7.2%?

use the following calculator.

https://arachnoid.com/finance/

15% down payment = .15 * 245,000 = 36,750.
subtract that from 245,000 and your mortgage loan needs to be 208,250.

for the first part, your inputs will be:

present value = 208,250
future value = 0
number of time periods = 30 * 12 = 360
interest rate per time period = 6.2/12 = .51666666666.....%
payments are to be made at the end of each time period.

click on PMT and the calculator tells you that the monthly payment will be
-1275.47

the present value is positive because it's money you receive.
the payment is negative because it's money you pay out.

your monthly payments will be 1275.47.

here's a display of what the calculator shows.

$$$

for the second part, everything will be the same except the interest rate per time period will be 7.2 / 12 = .6

click on PMT and the calculator tells you that the monthly payment will be -1413.58

the payment is negative because it's money you pay out.

here's a display of what the calculator shows.

$$$

you should not have to do this manually using formulas, but if you do, here's a reference on the most commonly used formulas.

if you try ehse formulas out, you should be able to duplicate the results from the use of this calculator, which is also referenced in the tutorial.

https://www.algebra.com/algebra/homework/Finance/THEO-2016-04-29.lesson#notes